# On additive error approximations to ##\##BQP

### v1

@@ 2.4 Additive Error Approximations
to a normalization $u(|x|)=2^{w(|x|)}$ with probability $1-\delta$ that runs in time polynomial in $1/\epsilon$ and $\log(1/\delta)$. The fact that this level of additive approximation is equal complexity for both #P and GapP follows immediately from another result of the paper, namely that additive approximations with a normalization $u$ are closed under subtraction.
Here we will study whether the complexity of analogous additive approximations to #BQP relations are of a different complexity. In order to do so, we adapt the definitions of additive approximations from Ref. [1] to relational problems. We also distinguish between the problem of returning an additive approximation to a relation and algorithms for doing so.
Definition 6 (Additive Error Approx.).
Given a relation $r\subset J\times\mathbb{Z}^{+}$ and a normalization $u:\mathbb{Z}^{+}\to\mathbb{R}^{+}$, the problem of returning an additive error approximation to $r$ with normalization $u$ to precision $\epsilon>0$ for any input $x\in J$, is to return a number $\tilde{r}(x,u)$ satisfying
@@
Intuitively this just means that the estimate needs to be close to at least one of the output $y$’s related to an input $x$. Using this definition we can then formally define the additive approximation algorithms we will be considering throughout this work.
Definition 7 (Additive Approx. Algorithm).
A classical (quantum) additive approximation algorithm for a relation $r\subset J\times\mathbb{Z}^{+}$ with normalization $u$ is a probabilistic (quantum) algorithm that given an input $x\in J$ returns a number $\tilde{r}(x,u)$ satisfying the additive error approximation requirements of Definition 6 with probability at least $1-\delta$ and runs in time (and size) polynomial in $|x|$, $1/\epsilon$ and $\log(1/\delta)$.
For the quantum algorithms, by polynomial time and size we mean polynomially many runs of polynomial-size quantum circuits and polynomial-time classical pre- and post-processing is allowed.
2.5 The Complexity of Additive Error Approximations to #BQP
@@
Here we present results related to additively approximating #BQP relations. The proofs for all of the theorems in this section are contained in Section 3.1. For our first result, we demonstrate that quantum additive approximation algorithms exist for all #BQP functions with normalization factors $u$ exponential in the number of qubits of the witness space.
Theorem 2.1.
For all relations $r\in$BQP with corresponding acceptance operators $\mathcal{V}_{x}^{r,|x|}$ on input $x$, there are quantum additive-error approximation algorithms with normalization factor $u(|x|)=2^{w_{r}(|x|)}$.
This result is analogous to the results in Ref. [1], but with the normalization now being determined by the number of qubits rather than the number of bits describing the witnesses. We next show that this level of additive approximation is in a sense the best possible.
Theorem 2.2.
@@
The problem of returning additive-error approximations for all #BQP relations $r$ with corresponding acceptance operators $\mathcal{V}_{x}^{r,|x|}$ on input $x$ to normalization $u(|x|)=2^{cw_{r}(|x|)}$ for positive $c<1$ satisfying $1-c=\Omega(\text{poly}(|x|)^{-1})$ is #BQP-hard.
Our next result implies that it is unlikely that there exists classical additive approximation algorithms for #BQP functions with the same normalization $u$ of Theorem 2.1.
Theorem 2.3.
Returning additive-error approximations with $u(|x|)=2^{w_{r}(|x|)}$ for all #BQP relations $r$ with corresponding acceptance operators $\mathcal{V}_{x}^{r,|x|}$ on input $x$ is BQP-hard.
This suggests that this level of approximation can not be achieved by any polynomial-time classical algorithm, and together with Theorem 2.1 suggests that this level of additive approximation to #BQP relations characterizes the difficulty polynomial time quantum computation.
@@
We next turn to looking for classical additive approximation algorithms to #BQP relations. Given the fact that exact counting classes are equivalent under polynomial-time reductions [5, 11], and that as shown in Ref. [1] there are efficient classical additive error approximations to all #P and GapP problems to some normalization $u$ there must exist classical additive approximations to #BQP relations. The key point that makes this consistent with the previous BQP-hardness result is that the reductions between the classes have a profound effect on the size of the witnesses, leading to a very different normalization $u$ between the quantum and classical approximation algorithms. We use the reduction in Ref. [5] that maps a #BQP relation to a GapP function with a witness size that results in a normalization $u$ which is significantly larger than the one in Theorem 2.1. A downside of the construction is that the achievable normalization likely depends on the gate set used.
Theorem 2.4.
@@
For a #BQP relation $r$ with input $x$, let $h_{r}\equiv h_{r}(|x|)$ be number of Hadamard gates in the circuit $V^{r,|x|}$ using the $\{\text{Toffoli},H,S\}$ gate set and let $t_{r}\equiv t_{r}(|x|)$ be the depth and $a_{r}\equiv a_{r}(|x|)$ be the number of ancilla qubits. Then classical additive error approximations exist with normalization factor $u(|x|)=2^{2t_{r}(w_{r}+a_{r})-(a_{r}+1)-h_{r}}$.
Note that this is in general not a particularly useful approximation, as $u$ is typically much greater than the range of values that the #BQP function can return in the first place.
2.6 The Relationship between Additive Error Approximations to #BQP and DQC1
@@
We now turn to analyzing the connection between additive approximations to #BQP relations and the complexity class DQC1. The proofs of the theorems in this section are presented in Section 3.2. This connection is motivated by a collection of results showing that additive approximations to certain problems related to or contained in #BQP, including the log-local quantum partition function and the log-local quantum density of states problems, are DQC1-complete [2, 9, 15]. Although the density of states problem is #BQP-complete, the reductions used affect the levels of additive error achievable, so conclusions are not immediate about the relationship between additive approximations to general #BQP relations and DQC1.
We now elucidate the connection between #BQP generally and DQC1 by showing that additive approximations to a subset of #BQP relations, namely those which have verifier circuits requiring only logarithmically many ancilla qubits, have the same computational power as DQC1. In order to show this we first need to define a version of #BQP with the notion of uniform generation that is used in DQC1.
Definition 8 (#BQP Alternative Version).
@@
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if for each input $x\in J$ there exists a polynomial-time generated quantum circuit $D_{x}^{r}$ and functions $c^{r}$,$s^{r}:\mathbb{N}\to[0,1]$ satisfying $c^{r}(n)-s^{r}(n)=\Omega(\text{poly}(n)^{-1})$ such that
$r(x)=[N_{\geq c^{r}(|x|)}(\mathcal{D}_{x}^{r})..N_{\geq s^{r}(|x|)}(\mathcal{D}_{x}^{r})].$
Our first result demonstrates that one only needs additive approximations to the subclass of #BQP relations with ancillas that scale logarithmically in the size of the number of witness qubits in order to solve DQC1 problems.
Theorem 2.5.
Returning additive-error approximations with $u(|x|)=2^{w_{r}(|x|)}$ for #BQP relations $r$ with corresponding verifier circuits $D_{x}^{r}$ with $a_{r}(|x|)=O(\log(w_{r}(|x|)))$ is DQC1-hard.
We next show that the ability to solve DQC1 decision problems a polynomial number of times allows for the computation of additive-error approximations to #BQP relations with logarithmically many ancillas.
Theorem 2.6.
The ability to solve DQC1 problems can be used to provide additive-error approximations with $u(|x|)=2^{w_{r}(|x|)}$ to all #BQP relations $r$ with corresponding verifier circuits $D_{x}^{r}$ on input $x$ such that $a_{r}(|x|)=O(\log(w_{r}(|x|)))$.
An immediate corollary of Theorems 2.5 and 2.6 is that returning additive approximations to #BQP relations with logarithmically many ancillas is complete for the alternative non-decision formulation of DQC1 described in Section 2.2.
3 Proofs
3.1 Additive Error Approximations to #BQP
In this section we provide proofs for Theorems 2.1, 2.2, 2.3 and 2.4.
Quantum Additive Approximation Algorithms to #BQP relations.
Here we prove Theorem 2.1.
@@ 3.1 Additive Error Approximations to #BQP
Given a verifier circuit $V^{r}_{x}$ corresponding to input $x$ acting on $w_{r}(|x|)$ witness qubits and $a_{r}(|x|)$ ancilla qubits with completeness and soundness parameters $c$ and $s$, we construct a new circuit $\tilde{V}^{r}_{x}$ with amplified completeness and soundness parameters which will be chosen later. Note that such a circuit can always be constructed while preserving the dimension of the witness subspace [19, 5]. The algorithm is then to generate $M=\text{poly}(|x|)$ uniformly random computational basis states for the witness subspace and send them through the new verifier circuit, and then count how many accept, and multiply this number by $2^{w_{r}(|x|)}/M$. Let $X_{i}$ be the indicator function that evaluates to $1$ if the $i$-th bit string accepts. Then this estimator is
@@
It can be shown that $\mathbb{E}[X]=\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]$ for the amplified version of the acceptance operator in Eq. (2), which due to the new $\tilde{c}$ and $\tilde{s}$ parameters satisfies
$N_{\geq c}(\mathcal{V})-(1-\tilde{c})2^{w_{r}(|x|)}\leq\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]\leq N_{\geq s}(\mathcal{V})+\tilde{s}2^{w_{r}(|x|)}\,.$ (8)
@@
$\textbf{Pr}(|X-\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]|\geq\epsilon 2^{w_{r}(|x|)})\leq$ $\frac{\text{Var}[X]}{\epsilon^{2}2^{2w_{r}(|x|)}}$ (10)
$=$ $\frac{1}{M\epsilon^{2}}\frac{\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]}{2^{w_{r}(|x|)}}\left(1-\frac{\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]}{2^{w_{r}(|x|)}}\right)$ (11)
$\leq$ $\frac{1}{M\epsilon^{2}}\,.$ (12)
@@
$N_{\geq c}(\mathcal{V})-(1-\tilde{c}+\epsilon)2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s}(\mathcal{V})+(\epsilon+\tilde{s})2^{w_{r}(|x|)}\,.$ (13)
@@
$N_{\geq c}(\mathcal{V})-\epsilon^{\prime}2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s}(\mathcal{V})+\epsilon^{\prime}2^{w_{r}(|x|)}\,.$ (14)
The success probability can then be amplified to $1-\delta^{\prime}$ with $\log(1/\delta^{\prime})$ applications of the previous procedure and then taking the median of the $\log(1/\delta^{\prime})$ estimates. We thus have a quantum algorithm that outputs an additive error estimate with normalization $u(|x|)=2^{w_{r}(|x|)}$ that has a total runtime and size polynomial in $|x|$, $1/\epsilon^{\prime}$ and $\log\left(1/\delta^{\prime}\right)$. ∎
Better approximation algorithms are #BQP-hard.
Next we prove Theorem 2.2 adapting a proof from Ref. [23].
@@
Given an input $x$ to a #BQP relation $r$ with corresponding verifier circuit $V_{x}^{r}$ and completeness and soundness parameters $c^{r}(|x|)$ and $s^{r}(|x|)$, define a new verifier circuit $V_{x}^{r^{\prime}}=V_{x}^{r}\otimes\mathbb{I}^{\otimes l}$ corresponding to another #BQP relation $r^{\prime}$ where the witness subspace has been expanded by $l$ qubits where $l$ will be chosen later. It then follows that $N_{\geq c^{r}(|x|)}(\mathcal{V}^{{}^{\prime}r}_{x})=2^{l}N_{\geq c^{r}(|x|)}(\mathcal{V}^{r}_{x})$ and $N_{\geq s^{r}(|x|)}(\mathcal{V}^{{}^{\prime}r}_{x})=2^{l}N_{\geq s^{r}(|x|)}(\mathcal{V}^{r}_{x})$. Now, assume an additive error estimate $\tilde{r}^{\prime}(x,u)$ to the #BQP instance corresponding to the circuit $V_{x}^{r^{\prime}}$ can be returned with a normalization $u_{r^{\prime}}(|x|)=2^{c(w_{r}(|x|)+l)}$ to some inverse polynomial precision $\epsilon<1$. Then the estimate $\tilde{r}=\tilde{r}^{\prime}(x,u)/2^{l}$ satisfies

### v2

@@ 2.5 Additive-Error Approximations
to a normalization $u(|x|)=2^{w(|x|)}$, with probability $1-\delta$, and that runs in time polynomial in $1/\epsilon$ and $\log(1/\delta)$. The fact that this level of additive approximation yields an equivalent complexity for both #P and GapP follows immediately from another result of the paper, namely that additive approximations with a normalization $u$ are closed under subtraction.
In this work we study whether the complexity of analogous additive approximations to #BQP relations are of a different complexity. In order to do so, we adapt the definitions of additive approximations from Ref. [1] to relational problems, as well as framing them as computational problems rather than algorithms.
Definition 5 (Additive Error Approximation).
Given a relation $r\subset J\times\mathbb{Z}^{+}$ and a normalization $u:\mathbb{Z}^{+}\to\mathbb{R}^{+}$, the problem of returning an additive-error approximation to $r$ with normalization $u$ to precision $\epsilon$, is to given an input $x\in J$, return a number $\tilde{r}(x,u)$ satisfying
@@
Intuitively this just means that the estimate needs to be close to at least one of the outputs $y$ related to an input $x$.
3 Summary of Results
3.1 The Complexity of Additive-Error Approximations to #BQP
@@
Here we present results related to additively approximating #BQP relations. The proofs for all of the theorems in this section are contained in Section 4.1. For our first result, we demonstrate that returning additive approximations to #BQP functions with normalization factors $u$ scaling exponentially in the number of qubits in the witness space is polynomial-time equivalent to BQP.
{thm}
[] Let $A$ be an oracle that given as input a #BQP relation $r$ with verifier circuits $V^{r,(n)}$, an input string $x$, and some $\epsilon=\Omega(\text{poly}(|x|)^{-1})>0$, returns an additive-error approximation to $r(x)$ with $u(|x|)=2^{w_{r}(|x|)}$. Then BQP $=$ BPPA.
The consequences of this theorem are two-fold. First, since the proof in Section 4.1 shows that the oracle $A$ can be simulated by a quantum subroutine, in analogy to the results in Ref. [1], there are efficient quantum algorithms for returning additive-approximations for all #BQP functions, but with the normalization $u$ now being determined by the number of qubits rather than the number of bits describing the witnesses. On the other hand, this result implies that returning additive approximations to #BQP functions to the stated $u$ is BQP-hard, so it is unlikely that they can be efficiently computed classically. We next show that this level of additive approximation is, in a sense, optimal.
{thm}
@@
[] The problem of returning additive-error approximations for all #BQP relations $r$ with corresponding acceptance operators $\mathcal{V}_{x}^{r}$ on input $x$ to normalization $u(|x|)=2^{cw_{r}(|x|)}$ for positive $c<1$ satisfying $1-c=\Omega(\text{poly}(|x|)^{-1})$ is #BQP-hard.
@@
Given the fact that exact counting classes are equivalent under polynomial-time reductions [5, 11], and as shown in Ref. [1] there are efficient classical additive-error approximations to all #P and GapP problems to some normalization $u$ there must exist classical additive approximations to #BQP relations. The key point that makes this consistent with the previous BQP-hardness result is that the reductions between the classes have a profound effect on the size of the witnesses, leading to a very different normalization $u$ between the quantum and classical approximation algorithms. We use the reduction in Ref. [5] that maps a #BQP relation to a GapP function with a witness size that results in a normalization $u$ which is significantly larger than the one in Theorem 3.1. A downside of the construction is that the achievable normalization likely depends on the gate set used.
{thm}
@@
[] For a #BQP relation $r$ with input $x$, let $h_{r}\equiv h_{r}(|x|)$ be number of Hadamard gates in the circuit $V^{r,(|x|)}$ using the $\{\text{Toffoli},H,S\}$ gate set and let $t_{r}\equiv t_{r}(|x|)$ be the circuit depth and $a_{r}\equiv a_{r}(|x|)$ be the number of ancilla qubits. Then classical additive-error approximations exist with normalization factor $u(|x|)=2^{2t_{r}(w_{r}+a_{r})-(a_{r}+1)-h_{r}}$. Note that this is in general not a particularly useful approximation, as $u$ is typically much greater than the range of values that the #BQP function can return in the first place.
3.2 The Relationship between Additive-Error Approximations to #BQP and DQC1
@@
Motivated by a collection of results showing that additive approximations to problems related to #BQP, including the log-local quantum partition function and the log-local quantum density of states problems, are DQC1-complete [2, 9, 15], we analyze the general connection between additive approximations to #BQP relations and the complexity class DQC1. Although the density of states problem is #BQP-complete, the reductions used affect the levels of additive error achievable, so conclusions are not immediate about the relationship between additive approximations to general #BQP relations and DQC1. We unify these results by showing that additive approximations to a subset of #BQP relations, those requiring only logarithmically many ancilla qubits, have the same computational power as DQC1. The proofs of this theorem is presented in Section 4.2.
{thm}
[] Let $A$ be an oracle that given as input a #BQP relation $r$ with verifier circuits $V^{r,(n)}$, where $a_{r}(n)=O(\log(w_{r}(n)))$, an input string $x$, and some $\epsilon=\Omega(\text{poly}(|x|)^{-1})>0$, returns an additive-error approximation to $r(x)$ with $u(|x|)=2^{w_{r}(|x|)}$. Then DQC1 $=$ BPPA.
Theorem 3.2 immediately implies the DQC1 completeness results of contained in [2, 15], as the verifier circuits for calculating the density of states only require logarithmically many ancillas, and polynomial-time reductions can be used to show that this result implies the DQC1 completeness of computing additive error approximations to partition functions in [9].
4 Proofs of Results
4.1 Additive-Error Approximations to #BQP
Prior to proving Theorem 3.1 we will introduce an alternative definition of #BQP and a BQP-complete problem that will make the reductions easier. The alternative version of #BQP below allows the promise gap to explicitly depend on the input.
Definition 6 (#BQP${}_{(c^{r},s^{r})}$).
@@ 4.1 Additive-Error Approximations to #BQP
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if there exists a uniformly-generated family of quantum circuits $V^{r,(n)}$ and a polynomial $q^{r}$ such that
$r(x)=[N_{\geq c^{r}(x)}(\mathcal{V}_{x}^{r})..N_{\geq s^{r}(x)}(\mathcal{V}_{x}^{r})]\,,$
where $c^{r},s^{r}:\{0,1\}^{*}\to[0,1]$ are efficiently-computable functions satisfying $c^{r}(x)-s^{r}(x)>1/q^{r}(|x|)$.
We show in Appendix 7 that Definition 6 is equivalent to Definition 2. The BQP-complete problem we introduce is that of estimating the average acceptance probability of verifier circuits. We prove its BQP-completeness in Appendix 8.
Definition 7 (Average Accept Probability of Verifier Circuits).
  1. 1.
    $\frac{\text{Tr}[\mathcal{V}_{x}]}{2^{w(|x|)}}\geq c(|x|)$ ,
  2. 2.
    or $\frac{\text{Tr}[\mathcal{V}_{x}]}{2^{w(|x|)}}\leq s(|x|)$ ,
promised one of these is the case, where $\mathcal{V}_{x}$ is the acceptance operator of Eq. (2) corresponding to $V^{(|x|)}$ on input $x$.
Using these two definitions, we are now ready to prove Theorem 3.1, restated here for convenience.
See 3.1
@@
In order to show BPPA $\subseteq$ BQP, we prove there exists a polynomial-time quantum subroutine which simulates $A$. Given a verifier circuit $V^{r,(|x|)}$ with completeness and soundness parameters $c$ and $s$, we construct a new circuit $\tilde{V}^{r,(|x|)}$ with amplified completeness and soundness parameters $\tilde{c}$ and $\tilde{s}$ to be chosen later. Note that such a circuit can always be constructed in polynomial time while preserving the dimension of the witness subspace [19, 5]. The algorithm is then to generate $M=\text{poly}(|x|)$ uniformly random computational basis states for the witness subspace and send them through the amplified verifier circuit, then count how many accept and multiply this number by $2^{w_{r}(|x|)}/M$. Let $X_{i}$ be the indicator function that evaluates to $1$ if the $i$th bit string accepts. Then the estimator is
@@
It can be shown that $\mathbb{E}[X]=\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]$ where $\tilde{\mathcal{V}}_{x}^{r}$ are the acceptance operators associated to the amplified verifier circuits $\tilde{V}^{r,(|x|)}$, as in Eq. (2). In terms of the amplified completeness and soundness parameters, we have
$N_{\geq c}(\mathcal{V}_{x}^{r})-(1-\tilde{c})2^{w_{r}(|x|)}\leq\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]\leq N_{\geq s}(\mathcal{V}_{x}^{r})+\tilde{s}2^{w_{r}(|x|)}\,.$ (8)
@@
$\text{Pr}(|X-\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]|\geq\epsilon 2^{w_{r}(|x|)})\leq$ $\frac{\text{Var}[X]}{\epsilon^{2}2^{2w_{r}(|x|)}}$ (10)
$=$ $\frac{1}{M\epsilon^{2}}\frac{\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]}{2^{w_{r}(|x|)}}\left(1-\frac{\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]}{2^{w_{r}(|x|)}}\right)$
$\leq$ $\frac{1}{M\epsilon^{2}}\,.$
@@
$N_{\geq c}(\mathcal{V}_{x}^{r})-(1-\tilde{c}+\epsilon)2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s}(\mathcal{V}_{x}^{r})+(\epsilon+\tilde{s})2^{w_{r}(|x|)}\,.$ (11)
@@
$N_{\geq c}(\mathcal{V}_{x}^{r})-\epsilon^{\prime}2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s}(\mathcal{V}_{x}^{r})+\epsilon^{\prime}2^{w_{r}(|x|)}\,.$ (12)
[…]
On the other hand, to show BQP $\subseteq$ BPPA, we use the alternative version of #BQP${}_{(c^{r},s^{r})}$ from Definition 6. Then querying $A$ for a #BQP${}_{(c^{r},s^{r})}$ relation $r$ on a polynomial number of inputs $x_{i}$ with normalizations $u_{i}=2^{w(|x_{i}|)}$, we show how this allows us to solve an instance of the Average Accept Probability of Verifier Circuits problem in Definition 7 with parameters $c(|x|)$ and $s(|x|)$, a BQP-complete problem (see Lemma 8). In particular, let $x\mapsto x_{i}=y_{i}x$ where $y_{i}$ is a bit string which encodes the correct completeness and soundness parameters to use for the corresponding input. Choosing a fixed length prefix-free encoding for $y_{i}$ ensures uniqueness of each concatenated bit string, and each $x_{i}$ has the same number of bits, which will be necessary to avoid repeated inputs. The completeness and soundness parameters, $c_{i}$ and $s_{i}$, are then determined by $y_{i}$.
Partition the interval $[0,1]$ into $M=\mathcal{O}(\text{poly}(|x|))$ subintervals $[s_{i},c_{i}]$, each of size $\Omega(\text{poly}(|x|)^{-1})$, corresponding to the different completeness and soundness parameters specified by $y_{i}$. Specifically, we consider the subintervals $[s_{i},c_{i}]$ where $c_{i}=(M-i)/M$ and $s_{i}=c_{i}-1/4M$ for $11\leq i\leq M-$ 1. Then querying $A$ for each problem input $x_{i}$ returns an estimate
$\hat{n}_{i}=N_{\geq c_{i}}+\delta_{i}+\varepsilon_{i}\,,$ (13)
with $\delta_{i}\in[0,N_{[s_{i},c_{i}]}]$ and $|\varepsilon_{i}|\in[0,\epsilon 2^{w(|x_{i}|)}]$. Letting $\hat{n}_{0}=0$, we define the two quantities $\hat{\chi}_{i}=\hat{n}_{i}-\hat{n}_{i-1}$ and $\Delta_{i}=c_{i}+1/2M$. Then we have
$=\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}+\sum_{i=1}^{M-1}(\delta_{i}+\varepsilon_{i})+\underbrace{\frac{1}{2M}(\delta_{M}+\epsilon_{M})}_{=0}\,,$
where the last term vanishes as a result of the fact that for the interval $[s_{M},c_{M}]$, we need not query the oracle and instead always return $2^{w(|x|)}$, meaning $\delta_{M},\epsilon_{M}=0$.
This quantity approximates the trace of the associated acceptance operator $\mathcal{V}_{x}^{r}$ of the Average Accept Probability of Verifier Circuits problem. In particular, note that
$|\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}_{[c_{i},c_{i-1}]}(\mathcal{V}_{x}^{r})|\leq\frac{1}{2M}N_{[c_{i},c_{i-1}]}\,,$ (15)
which then implies that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}(\mathcal{V}_{x}^{r})\bigg{|}\leq\frac{2^{w(|x|)}}{2M}\,.$ (16)
$\frac{1}{M}\sum_{i=1}^{M}\delta_{i}\leq\frac{2^{w(|x|)}}{M},\quad\text{and}\quad\frac{1}{M}\sum_{i=1}^{M}\varepsilon_{i}\leq\epsilon 2^{w(|x|)}\,,$ (17)
where the first inequality holds from the fact that $[s_{i},c_{i}]\cap[s_{j},c_{j}]=\emptyset$ for all $i\neq j$. In the second inequality we have used the fact that $2^{w(|x_{i}|)}=2^{w(|x|)}$ for all $i$, or rather, the size of the witness subspace for each input $x_{i}$ is the same size as that of $x$ even though the classical inputs may vary in size. Setting $\epsilon=1/M$, we add the errors from each contribution and find that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}\hat{\chi}_{i}-\text{Tr}(\mathcal{V}_{x}^{r})\bigg{|}\leq\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{2M}=\frac{5}{2}\frac{2^{w(|x|)}}{M}\,.$ (18)
Choosing $M>5/(c(|x|)-s(|x|))$ suffices to distinguish between the accepting and rejecting cases of Definition 7. Therefore, BQP $\subseteq$ BPPA. ∎
Next we prove Theorem 3.1, adapting a proof from Ref. [23], demonstrating that returning additive approximations to #BQP with better normalization factors is #BQP-hard.
See 3.1
@@
Given an input $x$ to a #BQP relation $r$ with corresponding verifier circuit $V^{r,(|x|)}$ and completeness and soundness parameters $c^{r}(|x|)$ and $s^{r}(|x|)$, define a new verifier circuit $V^{{}^{\prime}r,(|x|)}=V^{r,(|x|)}\otimes\mathbb{I}^{\otimes l}$ corresponding to another #BQP relation $r^{\prime}$ where the witness subspace has been expanded by $l$ qubits where $l$ will be chosen later. It then follows that $N_{\geq c^{r}(|x|)}(\mathcal{V}_{x}^{{}^{\prime}r})=2^{l}N_{\geq c^{r}(|x|)}(\mathcal{V}_{x}^{r})$ and $N_{\geq s^{r}(|x|)}(\mathcal{V}_{x}^{{}^{\prime}r})=2^{l}N_{\geq s^{r}(|x|)}(\mathcal{V}_{x}^{r})$. Now, assume an additive error estimate $\tilde{r}^{\prime}(x,u)$ to the #BQP instance corresponding to the circuit $V^{{}^{\prime}r,(|x|)}$ can be returned with a normalization $u_{r^{\prime}}(|x|)=2^{c(w_{r}(|x|)+l)}$ to some inverse polynomial precision $\epsilon<1$. Then the estimate $\tilde{r}=\tilde{r}^{\prime}(x,u)/2^{l}$ satisfies